# BlockPruner: Fine-grained Pruning for Large Language Models

### v2

@@ 4.1 Experimental Setups
In our main experiment, we utilize the Alpaca dataset (Taori et al., 2023) to calculate importance scores. For our method, we employ only 256 samples to compute perplexity, and we discuss the influence of varying sample sizes in Section 5.4. Moreover, we observe that ShortGPT and Relative Magnitude are not sensitive to different numbers of samples. The effect of sample size on ShortGPT and Relative Magnitude is detailed in Appendix E. Nevertheless, we used the same number of samples as our method for consistency. All experiments in this study are conducted using two RTX 4090 GPUs.
@@ 4.2 Main Results
Finally, we observe that while approaches such as ShortGPT and Relative Magnitude result in a significant decline in model performance across different tasks, BlockPruner stands out by avoiding such drastic reductions. This suggests that our proposed block pruning method effectively mitigates performance degradation during the pruning process. Due to space constraints, we have moved the details of pruning baselines and comparisons across various pruning ratios to Appendix F.
@@ 5.2 Redundancies Between MHA and MLP
The results in Figure 5 (left) for both models reveal a consistent tendency to initially remove only MHA blocks. As the pruning process progresses and more blocks are removed, the proportion of MHA blocks being pruned follows a zigzag downward trend. Notably, the curve for Llama2-13B shifts to the right compared to Llama2-7B, suggesting that the larger model contains more redundant MHA blocks. This is further emphasized in Figure 5 (right), where, at the same pruning ratio, Llama2-13B prunes more MHA blocks than Llama2-7B.
@@ Appendix C Details of Evaluations
Appendix D Post-training after Pruning
@@
Appendix E Sensitivity to Sample Size
@@
Appendix F Varying Pruning Ratios

### v3

@@ 4.1 Experimental Setups
In our main experiment, we utilize the Alpaca dataset (Taori et al., 2023) to calculate importance scores. For our method, we employ only 256 samples to compute perplexity, and we discuss the influence of varying sample sizes in Section 5.4. Moreover, we observe that ShortGPT and Relative Magnitude are not sensitive to different numbers of samples. The effect of sample size on ShortGPT and Relative Magnitude is detailed in Appendix H. Nevertheless, we used the same number of samples as our method for consistency. All experiments in this study are conducted using two RTX 4090 GPUs.
@@ 4.2 Main Results
Finally, we observe that while approaches such as ShortGPT and Relative Magnitude result in a significant decline in model performance across different tasks, BlockPruner stands out by avoiding such drastic reductions. This suggests that our proposed block pruning method effectively mitigates performance degradation during the pruning process. Due to space constraints, we have moved the details of pruning baselines and comparisons across various pruning ratios to Appendix I.
@@ 5.2 Redundancies Between MHA and MLP
The results in Figure 5 (left) for both models reveal a consistent tendency to initially remove only MHA blocks. As the pruning process progresses and more blocks are removed, the proportion of MHA blocks being pruned follows a zigzag downward trend. Notably, the curve for Llama2-13B shifts to the right compared to Llama2-7B, suggesting that the larger model contains more redundant MHA blocks. This is further emphasized in Figure 5 (right), where, at the same pruning ratio, Llama2-13B prunes more MHA blocks than Llama2-7B. Additionally, given that our pruning method tends to remove more MHA blocks at equivalent pruning ratios, it can significantly reduce the usage of the key-value (KV) cache (Pope et al., 2023) in MHA, which potentially accelerate the inference process. To validate this, we also conducted a comparison of the inference speed among various models obtained through different pruning methods, with the results detailed in Appendix E.
@@ Appendix C Details of Evaluations
Appendix D Ablation Experiments under Higher Sparsity
The motivation behind BlockPruner is to maintain model performance more effectively through fine-grained block pruning. Evaluating the performance differences of block pruning at varying granularities under higher sparsity is essential, as demonstrating effectiveness under such conditions would further corroborate our motivations and claims. In light of this, we conducted ablation experiments with higher sparsity ratios on LLaMA-2-7B and LLaMA-2-13B models and found that our motivation and claims remain valid. The results are presented in Table 4.
Model Unit Ratio(%) PPL($\downarrow$) Avg.Score
Llama2-7B Block 30.00/37.02/43.03 16.28/27.47/49.65 55.02/49.36/46.95
Layer 30.03/36.04/42.05 16.58/27.05/60.85 53.04/47.32/43.91
Llama2-13B Block 31.62/36.52/41.39 9.64/12.54/17.02 61.52/59.34/54.15
Layer 31.68/36.56/41.43 10.48/13.29/18.04 59.31/56.17/51.61
Table 4: Average score of BlockPruner at different pruning granularities under higher sparsity.
Appendix E Inference Speed after Pruning
Since our method reduces more MHA blocks at the same pruning ratio compared to other methods, it consequently reduces the usage of the KV cache at equivalent pruning ratios. This reduction can potentially accelerate the inference speed of LLMs. To verify that our method indeed accelerates the model’s inference process, we compared the time taken to generate 128 tokens by models obtained from different pruning methods with KV caches. To ensure the robustness of the experimental results, we repeated each configuration 20 times and took the average inference time as the result. As observed from the Table 5, our method achieves the highest speedup at comparable pruning ratios.
Model Method Ratio(%) Inference Time (ms) Speedup
Llama2-7B Original 0.00 4044.30 1.00
BlockPruner 24.00 2747.88 1.47
SliceGPT 24.47 3226.68 1.25
ShortGPT 24.03 3094.36 1.31
Llama2-13B Original 0.00 7285.73 1.00
BlockPruner 21.05 3873.20 1.88
SliceGPT 21.52 4099.08 1.78
ShortGPT 21.93 4111.65 1.77
Table 5: The inference speed differences among models obtained using different pruning methods, where “Original” denotes the unpruned model.
Appendix F Time Costs of Various Pruning Methods
Our approach relies on PPL to determine block importance, which requires calculating PPL before pruning, making it challenging to design a more efficient pruning strategy. We have compared other block importance metrics (in Section 5.3) but found that PPL still preserves the model’s performance best. Moreover, since our method better maintains model performance and pruning is one-time without increasing subsequent inference overhead, so we believe the trade-off is worthwhile. The comparison results of pruning times between BlockPruner and other methods are presented in Table 6.
Model BlockPruner SliceGPT ShortGPT RM LaCo
Llama2-7B 45 mins 2 hours 9 mins 2 mins < 1 mins 2 mins
Llama2-13B 2 hours 27 mins 3 hours 30 mins 2 mins < 1 mins 24 mins
Table 6: Execution time of BlockPruner and other pruning methods in the main experiment.
Appendix G Post-training after Pruning
@@
Appendix H Sensitivity to Sample Size
@@
Appendix I Varying Pruning Ratios